### v1

# Building Information Modelling MRR modular reconfigurable robot

Abstract
In situ robotic automation in construction is challenging due to constantly changing environments, a shortage of robotic experts, and a lack of standardized frameworks bridging robotics and construction practices. This work proposes a holistic framework for construction task specification, optimization of robot morphology, and mission execution using a mobile modular reconfigurable robot. Users can specify and monitor the desired robot behavior through a graphical interface. Our framework identifies an optimized robot morphology and enables automatic real-world execution by integrating Building Information Modelling (BIM). By leveraging modular robot components, we ensure seamless and fast adaption to the specific demands of the construction task. Experimental validation demonstrates that our approach robustly enables the autonomous execution of robotic drilling.
Note to Practitioners
This work addresses the challenge of automating robotic drilling on construction sites and introduces a general framework that also supports other construction tasks. Construction sites are unpredictable and constantly changing, making automation difficult. To tackle this, we propose a modular robotic system that can adapt to specific tasks. We use Building Information Modelling (BIM) data to automatically determine the best robot configuration by optimizing factors such as setup time and robustness to positioning errors. Through an intuitive interface, workers can specify tasks, after which they assemble the robot following the provided instructions.
This approach enables the same hardware to be deployed across various scenarios while accounting for site-specific challenges. The core innovation lies in combining modular robotics with automated optimization, task planning, and motion planning to minimize the need for user intervention while delivering robots specifically tailored to each task. To achieve this, the work extends existing methods for modular robot design, enabling consideration of multiple, potentially conflicting objectives in the optimization process. Experimental results demonstrate the ability of the system to perform autonomous drilling with high precision. While the system requires initial calibration, future improvements in vision and force control could further enhance its accuracy and robustness. Beyond drilling, this framework has the potential to support a wide range of applications, such as welding, painting, and assembly, offering a practical and adaptable solution for diverse construction tasks.
Index Terms:
Robotic Driling, Modular Robots, BIM
I Introduction
Figure 1: Section A shows the available robot modules for the manipulator, composed of six joints (green), three static links (silver), and a custom end-effector. Section B shows the assembled robotic arm on a mobile base and the experimental setup.
In recent decades, robotic automation in manufacturing significantly increased productivity [1]. To this day, a comparative technology leap in the construction sector has not occurred [2, 3, 4, 5]. Additionally, harsh, non-ergonomic working conditions result in labor shortages in the construction industry [6]. Robotic automation in construction, however, requires versatility, adaptability, and robustness as construction sites are complex, constantly changing, potentially hazardous, and hard to navigate [7].
The main contribution of this paper is a framework for in situ robotic automation that features (i) an autonomous decomposition of high-level instructions (e.g., “drill holes” or “paint wall”) into executable skills, (ii) identification of an optimized module composition for the robot arm using BIM data of the environment, (iii) online motion planning, and (iv) adaptions to positioning errors of the mobile base. The only actions required by the human worker are specification of the task and assembly of robot modules once the optimization of their composition terminates. While our framework is not limited to a particular task, we focus on autonomous drilling as shown in  Figure 1. A supplementary video of the real-world experiment and more can be found on the project website modular-construction-robot.cps.cit.tum.de.
II Related Work
This section discusses previous applications of mobile manipulators in construction and, specifically, drilling use cases. Further, we provide an overview of current methods for optimizing modular robot compositions.
II-A Robotic automation in building construction
Some construction tasks, such as cutting bricks or insulation panels, can be automated by special machinery or onsite manufacturing processes [8]. However, highly accurate and flexible interactions with the environment necessitate mobile robotic systems that can self-localize [9]. Navigating the highly unstructured environments of construction sites challenges the sensing and perception capabilities of robots. In [10], 3D scans taken with onboard sensors were used to enable navigation and localization of a bricklaying robot in situ. The perception can be further enriched with data from as-built BIM [11].
Leveraging BIM data in construction has been shown to increase worker productivity [12] and accelerate the deployment of construction robotics [13]. However, as reported by [14], there is still a lack of research on integrating BIM with onsite robotic systems. Previous work proposed the integration of BIM data into a mission parametrization toolchain for non-experts and validated its effectiveness through a paint-spraying experiment [15].
Particularly challenging use cases in construction robotics are precise power tool works on walls or ceilings, such as grinding or drilling holes. The challenges stem from uncertainties in robot perception and disturbances induced by operating the power tools. Recent works present an aerial system for drilling holes limited to free space areas and vertical walls or facades [16] or the development of a mobile manipulator with predefined reach and dexterity [17].
Robots with a fixed kinematic, as presented by industrial manufacturers111for instance, “Hilti Jaibot”, “Baubot”, or “Schindler R.I.S.E” or in previous research [18, 19, 20], may not provide the necessary flexibility required in construction [21]. Reconfigurable modular robots – comprising modular joints, links, and sensor modules – offer better transportability to and on the construction site and can also be adapted to specific requirements in situ. This aligns with previous research on digital in situ fabrication [22], which highlights that versatile, modular, and reusable solutions and real-world experiments are required to enable innovation in construction automation.
II-B Optimizing modular robot composition
The optimization of robot morphologies, and more specifically, finding an optimal module composition for reconfigurable robots, has attracted attention for decades [23, 24, 25, 26, 27]. Some approaches introduce heuristics and surrogate performance criteria to enable a brute-force search over reduced search spaces [28, 29, 30]. Most approaches, however, use discrete black-box optimization algorithms [31, 32, 33]. Since the early work of Karl Sims [34], genetic algorithms (GAs) have been a common choice for the optimization of modular robot morphologies. The work in [35] deploys a genetic algorithm (GA) to co-optimize the morphology and base placement of a modular robot. A lexicographic genetic approach is proposed in [36] to increase the efficiency and interpretability of the evolutionary optimization process. While our work builds upon [36], we extend it to multi-objective problems to explicitly model the tradeoff between performance and robustness. Further, in contrast to previous work, we perform an experimental validation of the optimized designs.
II-C Robustness and multi-objective optimization
It is well-known that a purely simulation-based optimization of real-world systems can lead to suboptimal performance during deployment – a phenomenon frequently termed the reality gap [37]. This is particularly problematic in unstructured or partially structured environments, like construction sites, where conditions are highly dynamic and unpredictable, and simulation often fails to capture the full complexity of real-world variability. Especially in learning-based approaches, domain randomization has been successfully applied to tackle this issue [38, 39, 40]. However, past works indicate a tradeoff between the specificity and generality of a solution [41, 42]. In this work, we explicitly model this tradeoff by formulating a multi-objective optimization problem with performance and robustness against positioning errors as competing objectives to find Pareto-optimal solutions. To this end, we incorporate the well-known nondominated sorting-based genetic algorithm (NSGA-II[43] into the lexicographic optimization procedure previously introduced in [36].
III Framework
Figure 2: The structure of the proposed framework from optimization of a robot model to mission execution. During mission preparation, inputs from the user interface are used to perform robot synthesis. The mission execution and robot control block run during deployment.
Our framework is shown in Figure 2. It can be separated into four modules: The user interface and robot synthesis modules are required for mission preparation, whereas mission execution and robot control run during deployment. A user generates and parameterizes a mission through the BIM-based user interface presented in [15]. Based on the user input, the robot synthesis module searches for the optimal configuration of the robot modules in simulation to execute the provided mission and proposes it to the user. Next, the user assembles the modular robot and activates its control setup. Once the robot is ready, the initially configured mission is sent to the robot. Finally, the robot control module adapts an initially planned trajectory based on visual calibration and executes it on the real robot.
III-A Robot Synthesis
Reconfigurable modular robot compositions are simulated and optimized with Timor [44]. We rely on the modular robot modeling language introduced in CoBRA [45] and a custom parser module (CoBRA BT Parser) to integrate the user interface and the BIM model with Timor. After the optimization algorithm described in Section IV identifies an optimized module composition and a motion trajectory, the worker can proceed with the physical robot assembly. The self-identification of the robot [46] or automatic model generation [44] can be used to obtain a kinematic and dynamic model of the robot once its modules are provided. The motion trajectory is stored and later passed to the robot control module.
III-B Robot Control
The robot control follows a modular setup. While our framework is agnostic to the specific software being used for control, we integrated the following components for deployment:
  • The mobile base navigation, including local and global path planning, localization, and omnidirectional control, is based on the navigation2 framework [47, 48].
  • The manipulator control, including collision-free trajectory planning and execution, is based on Timor and joint impedance control using the manipulator control library [^2] https://github.com/FraunhoferItalia/fhi_manipulator_control_stack. .
  • The calibration module based on OpenCV [^3] https://github.com/opencv/opencv includes a pipeline that detects ArUco landmark poses [49] with a camera to calibrate the robot position with respect to the environment. As the mobile base cannot move with arbitrary precision, even after calibration, the manipulator might not exactly be in the desired position – in this paper, we refer to this remaining error as the base positioning error.
IV Modular Robot Composition Optimization
The following section introduces our method for optimizing modular robot compositions given a task description and assumptions about the positioning accuracy of the mobile base. Throughout this section, we follow the notation shown in Table II.
IV-A Preliminaries
TABLE I: Scores for exemplary robots and multiple objectives.
${\mathbf{r}}_{1}$ ${\mathbf{r}}_{2}$ ${\mathbf{r}}_{3}$ ${\mathbf{r}}_{4}$
$f_{fast}$ 1 2 2 3
$f_{cheap}$ 2 1 3 1
$f_{robust}$ 3 2 2 2
As a running example, consider a ranking of robots $\{1,2,3,4\}$, based on how fast they can perform a specific task ($f_{fast}$), how cheap the corresponding hardware is ($f_{cheap}$), and how robust they are to external disturbances ($f_{robust}$). Table I shows the scores ${\mathbf{r}}_{i}$ of four exemplary robots. The higher the score, the better the robot. There are multiple objectives, so no clear “best” robot exists. In the following sections, we will introduce a ranking-based selection method that considers multiple objectives.
Partially Ordered Set. A partially ordered set is a tuple $({\mathbb{A}},\leq)$ of a set ${\mathbb{A}}$ and the order relation $\leq$, which is reflexive, antisymmetric, and transitive [50, Def. 1.1].
Lexicographic order. Let ${\mathbf{a}},{\mathbf{b}}$ with $\mathrm{len}({\mathbf{a}})=\mathrm{len}({\mathbf{b}})=n$ be two sequences with elements $a_{i},b_{i}\in({\mathbb{S}},\leq)$. We define the lexicographic order of ${\mathbf{a}}$ and ${\mathbf{b}}$ as [51, Def. 1.8]
$\begin{split}{\mathbf{a}}>{\mathbf{b}}\Leftrightarrow&\exists\,i\in\left[n\right]:{\textnormal{a}}_{i}>{\textnormal{b}}_{i}\land\forall j<i:{\textnormal{a}}_{j}={\textnormal{b}}_{j}\,,\\ {\mathbf{a}}={\mathbf{b}}\Leftrightarrow&\forall\,i\in\left[n\right]:{\textnormal{a}}_{i}={\textnormal{b}}_{i}\,.\end{split}$ (1)
Intuitively, this defines a hierarchical order over the elements of every sequence by decreasing importance.
Let $(f_{fast},f_{cheap},f_{robust})$ be the sequence of lexicographic fitness values for the exemplary robots. As the fourth robot is the fastest one and robot three is cheaper than robot two, while they are equally fast, we obtain the lexicographic order ${\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}>{\mathbf{r}}_{1}$. However, if we chose our fitness sequence to be $(f_{robust},f_{fast},f_{cheap})$, the order would be ${\mathbf{r}}_{1}>{\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}$.
Pareto optimality. Let ${\mathbf{f}},{\mathbf{g}}$ with $\mathrm{len}({\mathbf{f}})=\mathrm{len}({\mathbf{g}})=n$ be two sequences with ${\textnormal{f}}_{i},{\textnormal{g}}_{i}\in({\mathbb{F}},\leq)$. We say “${\mathbf{f}}$ dominates ${\mathbf{g}}$” if it contains a higher value for at least one objective and if all other objectives are at least equal. Formally,
${\mathbf{f}}{\succ}{\mathbf{g}}\Leftrightarrow\exists\,i\in\left[n\right]:{\textnormal{f}}_{i}>{\textnormal{g}}_{i}\land\forall\,i\in\left[n\right]:{\textnormal{f}}_{i}\geq{\textnormal{g}}_{i}\,.$ (2)
An element ${\mathbf{f}}$ is Pareto-optimal with respect to a set ${\mathbb{F}}$ if it is not dominated by any element in this set [52, eq. 4.59].
Let $(f_{fast},f_{cheap},f_{robust})$ be a multi-objective fitness sequence for the exemplary robots. We can see that ${\mathbf{r}}_{3}{\succ}{\mathbf{r}}_{2}$ as both robots are equally fast and robust, but robot three is cheaper. However, robots one, three, and four are all Pareto-optimal as none of them is dominated by any other robot.
Genetic algorithm. Genetic algorithms change the population of different solution candidates (individuals) to maximize an objective, called the fitness function [53]. Every individual is represented by a genome that encodes its unique attributes. For a fixed number of iterations, a new population is created by selecting a limited number of individuals from the previous population (selection), creating offsprings by combining attributes of existing individuals (crossover), and mutating some of the genomes (mutation). A cycle of selection, crossover, and mutation is called a generation.
NSGA-II. We write $\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ and $\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ whenever we compute the nondomination rank or the crowding distance of an individual ${\mathbf{a}}$ with respect to a population ${\mathbb{A}}$ according to the NSGA-II algorithm as proposed in [43].  -A provides an informal introduction to these metrics.
IV-B Lexicographic nondominant sorting optimization
TABLE II: Nomenclature
General notation:
${\mathbf{a}},{\mathbf{b}},{\mathbf{c}}$ Sequences ${\mathbb{A}},{\mathbb{B}},{\mathbb{C}}$ Sets
${\bm{a}},{\bm{b}},{\bm{c}}$ Vectors ${\bm{A}},{\bm{B}},{\bm{C}}$ Matrices
$\alpha,\beta,\gamma\in\mathbb{R}$ Scalar variables ${\mathcal{A}},{\mathcal{B}},{\mathcal{C}}$ Multisets
$\left[n\right]$ The set $\{1,2,\dots,n\}$ $({\mathbb{A}},\leq)$ Partially ordered sets
$\mathrm{Tr}({\bm{M}})$ The trace of a matrix $|{\mathbb{A}}|$ The cardinality of a set
Important symbols:
${\bm{b}}$ Pose $(x,y,\theta)$ of the mobile base
${\mathcal{M}}$ Available modules
${\mathbf{m}}$ A sequence of modules, defining a robot
${\bm{q}}$ Joint angles of the manipulator
${\mathbb{Q}}$ Robot joint space
${\mathbb{O}}$ Environment obstacles
${\mathbf{g}}=({\mathbf{p}},t)$ A goal with desired pose ${\mathbf{p}}$ and tolerance $t$
${\mathbf{p}}=({\bm{t}},{\bm{R}})$ A pose with translation ${\bm{t}}\in\mathbb{R}^{3}$ and orientation ${\bm{R}}\in SO(3)$
${\mathbb{O}}_{{\mathbf{m}},i}\left({\bm{q}}\right)$ Space occupied by the $i$th link of ${\mathbf{m}}$ for joint angles ${\bm{q}}$
Custom Operators:
$\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ The crowding distance of a multi-objective value ${\mathbf{a}}$ with respect to its nondomination front in set ${\mathbb{A}}$
$\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ The nondomination rank of a multi-objective value ${\mathbf{a}}$ with respect to the set ${\mathbb{A}}$
$\mathrm{count}\left(x,\cdot\right)$ The number of elements $x$ within a sequence ($\mathrm{count}\left(x,{\mathbf{x}}\right)$) or multiset ($\mathrm{count}\left(x,{\mathcal{X}}\right)$)
$\mathrm{len}({\mathbf{a}})$ The number of elements in a sequence
$\mathrm{size}(m)$ The size of a robot module, i.e., the Euclidean distance between its proximal and distal connection interfaces
We perform a mixed Pareto-lexicographic optimization based on the approach proposed in [36]. This work introduces a genetic algorithm with a selection that is based on a purely lexicographic fitness function. Notably, the approach initially targets a single-objective optimization. The introduction of a lexicographic fitness function is motivated by computational efficiency: By introducing multiple surrogate objectives, the selection process of the genetic optimization can be performed without evaluating the computationally expensive primary objective for a majority of solution candidates.
Our work combines this approach with the NSGA-II algorithm [43] to achieve computationally efficient multi-objective optimization. We compute two sequences ${\mathbf{c}}_{i}$ and ${\mathbf{x}}_{i}$ for every individual $i$ in a population. The sequence ${\mathbf{c}}_{i}$ contains $m$ values that measure compliance with $m$ constraints that can be easily evaluated, such as the sufficient size of a robot or the absence of self-collisions. We define these constraints so that the $j$th value of ${\mathbf{c}}_{i}$, $c_{i,j}$ is 00, if solution $i$ satisfies constraint $j$, and negative, otherwise. In addition to previous work, we introduce the sequence
${\mathbf{x}}_{i}$ =$=\begin{cases}(a_{i},b_{i},\dots)&\text{, if }{\mathbf{c}}_{i}=\bm{0}\\ (-\infty,-\infty,\dots)&\text{, otherwise,}\end{cases}$ (3)
where $a_{i},b_{i},\dots$ can be arbitrary objectives, such as the monetary cost or robustness of an individual solution. We write ${\mathbb{X}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}$ and define the lexicographic fitness as the sequence
${\mathbf{f}}_{i}=(\underbrace{{\textnormal{c}}_{i,1},{\textnormal{c}}_{i,2},\dots,{\textnormal{c}}_{i,m}}_{\text{constraint satisfaction}},\underbrace{\mathrm{ndr}({\mathbf{x}}_{i},{\mathbb{X}})}_{\text{nondomination rank}},\underbrace{\mathrm{cd}\left({\mathbf{x}}_{i},{\mathbb{X}}\right)}_{\text{crowding distance}})\,.$ (4)
By performing a lexicographic selection based on this fitness, we can efficiently rule out candidates that do not fulfill the constraints. In addition, the nondomniation rank fitness from  (4) allows us to perform multi-objective selection under consideration of ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}$. Instead of a single solution, our optimization results in a set of Pareto-optimal solutions from which an operator can pick.
IV-C Task representation
We obtain a formalized task representation by parsing the BIM model and user input as described in Sec. III. The representation of goals, obstacles, tolerances, and robots follows the modeling language introduced by the CoBRA benchmark [45]. A formal definition for our specific use case is included in -B and -C.
IV-D Fitness functions
For motion planning, we use the toolbox for industrial modular robotics (Timor) [44]. We obtain base movement ${\bm{b}}(t)$ and joint angles ${\bm{q}}\left(t\right)$ over time from a black-box path planner
$\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},{\mathbb{C}},t\right)=\begin{bmatrix}{\bm{b}}(t)\\ {\bm{q}}\left(t\right)\end{bmatrix}\,.$ (5)
The planner guarantees satisfaction for all constraints ${\mathbb{C}}$ considered in this work, such as collision avoidance. However, inverse kinematics solutions to reach all goals ${\mathbb{G}}$ might not always exist. Further, as we rely on numeric optimization, we might not be able to find them with perfect accuracy. To obtain a scalar expression of the errors ${\mathbf{e}}{}=\left({\bm{t}_{e}},{\bm{R}_{e}}\right)$ (21) given by an imperfect solution, we introduce weighting factors $w_{t},w_{R}$ and error measures $d_{t}:\mathbb{R}^{3}\to\mathbb{R}$ and $d_{R}:SO(3)\to\mathbb{R}$ and define the distance function
$d({\mathbf{e}})=w_{t}d_{t}({\bm{t}_{e}})+w_{R}d_{R}({\bm{R}_{e}})\,.$ (6)
While the measures can be chosen arbitrarily, we usually use the Euclidean norm for $d_{t}$ and define $d_{R}$ as the Euler angle of rotation (22). End effector poses can easily be computed using the forward kinematics $FK:\mathbb{R}^{3}\times{\mathbb{Q}}\to SE(3)$. This allows us to obtain the minimum pose error for goal ${\mathbf{g}}_{i}$:
${\mathbf{e}}^{*}_{i}(\bm{\xi}\left(t\right),{\mathbf{p}}_{i})$ $={\mathbf{e}}{({\mathbf{p}}_{i},FK(\bm{\xi}\left(\hat{t}\right)))}\text{, where}$ (7)
$\hat{t}$ $=\operatorname*{arg\,min}_{t}d({\mathbf{e}}({\mathbf{p}}_{i},FK(\bm{\xi}\left(t\right))))\,.$
For better readability, we omit the arguments when they are clear from context and write ${\mathbf{e}}_{i}^{*}$. If the goal error is within the goal tolerance for every goal in a task, we say that the path $\bm{\xi}\left(t\right)$ is feasible.
Let ${\mathbb{O}}_{{\mathbf{m}},i}\left({\bm{q}}\right)$ be the space that is occupied by the $i$th link of robot ${\mathbf{m}}$ with $n$ joints and joint configuration ${\bm{q}}$. The following constraints for joint limits (8), self-collisions (9), and collisions with the environment (10) all evaluate to true or false:
$c_{\text{lim}})$ $\forall t\in[T]:{\bm{q}}\left(t\right)\in{\mathbb{Q}}$ (8)
$c_{\text{sc}})$ $\forall\left(i,t\right)\in\left[n\right]\times[T],\,\forall j<i:$
$\qquad\qquad{\mathbb{O}}_{{\mathbf{m}},i}\left({\bm{q}}\left(t\right)\right)\cap{\mathbb{O}}_{{\mathbf{m}},j}\left({\bm{q}}\left(t\right)\right)=\emptyset$ (9)
$c_{\text{col}})$ $\forall\left(i,t\right)\in\left[n\right]\times[T]:{\mathbb{O}}\cap{\mathbb{O}}_{{\mathbf{m}},i}\left(\bm{\xi}\left(t\right)\right)=\emptyset\,.$ (10)
We define the fitness functions for constraint satisfaction of robot ${\mathbf{m}}$ as follows.
${c_{1}}$ $=\min\left(0,\sum_{i=1}^{\mathrm{len}({\mathbf{m}})}\bigg{(}\mathrm{size}({\textnormal{m}}_{i})\bigg{)}-\max_{i\in[|{\mathbb{G}}|]}\left(\left\lVert{\bm{t}}_{i}\right\rVert_{2}\right)\right)$ (11)
${c_{2}}$ $=\min\left(0,\min_{i}\left(\mathrm{count}\left({\mathcal{M}},{\textnormal{m}}_{i}\right)-\mathrm{count}\left({\mathbf{m}},{\textnormal{m}}_{i}\right)\right)\right)$ (12)
${c_{3}}$ $=-\sum_{t>T_{cal}}\left\lVert{\bm{b}}(t)-{\bm{b}}(T_{cal})\right\rVert$ (13)
${c_{4}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}}\},{\mathbb{G}},t\right))\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (14)
${c_{5}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}}\},{\mathbb{G}}\right),t)\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (15)
${c_{6}}$ =$=\begin{cases}0&\text{, if }\mathrm{feasible}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},\{c_{\text{lim}},c_{\text{sc}},c_{\text{col}}\},{\mathbb{G}},t\right))\\ d({\mathbf{e}}^{*}_{i})&\text{, otherwise.}\end{cases}$ (16)
The functions in (11) and (12) can be evaluated without obtaining a kinematic or dynamic model of the robot and indicate whether the individual modules of the robot are large enough to reach all goals and if they are available in situ, respectively. The constraint in (13) forces the base not to move after the visual calibration is performed ($t=T_{cal}$, see also Figure 4). The remaining functions successively evaluate whether the joint path is valid regarding the joint limits (14), self-collisions (15), and collisions with the environment (16).
If all constraints are satisfied, we compute the values for our multi-objective fitness criteria. The compactness
$f_{c}$ $=-\sum_{i=1}^{n}\mathrm{size}(m_{i})$ (17)
indicates the negative value of the total size of all modules combined. High compactness usually allows faster transportation and a smaller lever for the forces induced by drilling.
The robustness fitness function indicates the maximum positioning error our motion planner can account for online. After visual calibration, we obtain a placement error $\Delta_{b}=(\Delta_{x},\Delta_{y},\Delta_{\theta})^{T}$ between the desired and the actual base placement of the robot. The adjustment function $\mathrm{replan}(\bm{\xi}\left(t\right),\Delta_{b})$ indicates whether there is an adjustment $\bm{\varepsilon}$ that, when applied to the joint path ${\bm{q}}\left(t\right)$, results in the end effector of the robot still following the desired motion. Finally, we define the robustness of a robot ${\mathbf{m}}$ with respect to positioning errors of the mobile base as
$f_{r}=$ $\max\{\delta\in[0,1]:\mathrm{replan}(\bm{\xi}\left({\mathbf{m}},{\mathbb{G}},{\mathbb{C}},t\right),\delta\Delta_{max})\}\,.$ (18)
The maximum displacement $\Delta_{max}$ can be chosen based on the available hardware and prior experience. We choose $\begin{bmatrix}20\text{cm}&20\text{cm}&15\deg\end{bmatrix}^{T}$ for our experiments. The above definition does not discriminate between robustness against positioning errors in the $xy$-plane and orientation errors $\Delta_{\theta}$. If necessary, it can easily be extended to multiple robustness scores, one for every possible error. In practice, we experienced that positioning errors were mainly caused by slipping wheels, rendering an equal consideration of the different error modes plausible.
Lastly, the reconfiguration time fitness indicates the time necessary to reconfigure the current module composition ${\mathbf{m}}$ to match a new, desired configuration ${\mathbf{m}}_{d}$. If the current composition is unknown or non-existent, for example, because the modules are currently disassembled, we set ${\mathbf{m}}$ to the empty tuple, so the reconfiguration fitness measures the initial time for setup. Assume that ${\mathbf{m}}$ and ${\mathbf{m}}_{d}$ share $n_{c}$ common modules at the beginning of their kinematic chain. Then, the reconfiguration time fitness is given by
$f_{t}$ $=-t_{m}\left(\mathrm{len}({\mathbf{m}})+\mathrm{len}({\mathbf{m}}_{d})-2n_{c}\right)\,.$ (19)
Here, we assume that removing or adding a module takes a time of $t_{m}$, regardless of the module.
V Experiments
We conducted experiments on a drilling use case in the Fraunhofer Italia ARENA at the NOI Techpark in Bolzano, Italy. The experimental setup is shown in Figure 1. It consists of a movable wall with four blocks (brick, sandstone brick, wood, and insulation material) that is modeled in the used BIM file [^4] https://github.com/FraunhoferItalia/rosbim/tree/humble-dev/rosbim_example_models.
V-A Robot modules
We use the robot modules introduced in [46] and shown in Figure 1A, together with an RB-VOGUI mobile base555 https://robotnik.eu/products/mobile-robots/rb-vogui-en/ with a custom cabinet top setup and two lidars (SICK expert S300). In total, we use one mounting base, three elbow- and straight joint modules each, three passive links666 https://github.com/FraunhoferItalia/fhi_resources with a length of $100$, $200$, and $500$ millimeters, respectively, as well as a custom-made end effector, composed of a drill tool (Einhell Herocco 18/20 with custom 3D printed support) and an RGB-D camera (Intel Realsense D435i).
V-B Validation in simulation
(a) Normalized compactness and robustness of Pareto-optimal results within 100 individual runs. The heights of the boxes indicate the relative frequency of a particular fitness score observed. The blue line indicates the Pareto optimal results concerning all trials.
(b) The best robustness score that could be achieved with a limited reconfiguration time.
Figure 3: Aggregated results for 100 different random initial populations. Shaded areas indicate 95% confidence intervals using bootstrapping.
[…]
Figure 3a shows the Pareto-optimal solutions of each trial regarding compactness and robustness. The compactness is normalized for better readability, i.e., 00 indicates the largest and $1$ the most compact robot present in the final solution set, respectively. As our optimization is not complete, the theoretic Pareto front remains unknown. Instead, we report the evidence-based Pareto front by aggregating over all trials in blue. Further, Figure 3b shows the tradeoff between robustness and reconfiguration time, starting from an exemplary initial configuration.
V-C Real-world experiments
Figure 4: The simplified behavior tree of the drilling experiment on a sandstone brick with the six-degrees-of-freedom configuration of the modular robotic arm.
Figure 4 shows a simplified version of the behavior tree that consists of one sequence of robot actions used in the experiments. The full tree additionally includes fallback nodes for error handling. Pictures of the drilled holes can be seen in the most-right snapshot in Figure 4 and in Figure 6. A full video of the mission execution is available on the project website modular-construction-robot.cps.cit.tum.de.
Figure 5: Among the conducted experiments, we measure absolute errors between $6$ cm and $16$ cm regarding the base position and between $0.7^{\circ}$ and $12.6^{\circ}$ regarding its orientation. The figure shows the positioning errors of the base with respect to the ArUco markers and the maximum joint angle adjustment during execution of the drilling task.
Based on prior experience with the mobile base, we chose a minimum robustness score of $\delta=0.8$. Consecutively, we selected two solutions from our optimization; one with a six-degrees-of-freedom (DoF) robot arm and one with a five-DoF robot arm. For every configuration, we performed three drilling missions on the sandstone brick.
Figure 5 shows the absolute error of the mobile base position relative to the nominal drilling position for each experiment, expressed in terms of $e_{x}$, $e_{y}$, and $\theta$. Moreover, it indicates the adjustment necessary to correct the initial motion plan according to the positioning error. We compute this score as the maximum change in a single joint angle over all joint configurations in the drilling path $\bm{\xi}\left(t\right)$, i.e.,
$\hat{\varepsilon}=\max_{t\in[T]}\left(\left\lVert{\bm{q}}(t)-\hat{{\bm{q}}}(t)\right\rVert_{\infty}\right)\,,$ (20)
where ${\bm{q}}$ is the initially planned joint configuration, and $\hat{{\bm{q}}}$ is the one observed during execution. We chose this metric as it can easily be compared between robots with varying DoF; however, the overall trend shown in Figure 5 persists over different choices of norms.
The first experiment was conducted with the six-DoF manipulator. Figure 6 shows the sandstone brick after drilling three holes in a horizontal line. The relative vertical precision was within one centimeter.
Figure 6: The sandstone bricks after drilling three holes in a vertical line. The repeatability precision of the whole approach, including navigation and calibration is within $1$ cm.
Initially, the modules were disassembled, as depicted in Figure 1A. Upon receiving the configuration for the robot (a six-DoF setup consisting of the following module sequence: straight, elbow, passive 100mm, straight, elbow, straight, elbow, end effector), the operator assembled the robot in ten minutes. For the second experiment, the robot consisted of the following module sequence: straight, elbow, passive 100mm, straight, elbow, elbow, end effector. The time required to reconfigure the robot and launch the mission was reduced to about five minutes, as the first five modules were shared with the previous configuration.
V-D Discussion
A compactness-robustness tradeoff can clearly be seen in both mean and Pareto-optimal scores. The results shown in Figure 3b indicate a similar tradeoff concerning the reconfiguration time. When starting from an initially assembled robot from a previous task, the first feasible robot for drilling was identified within three minutes. However, if high robustness is essential, a reconfiguration time of over 20 minutes is necessary to assemble the corresponding robot [^7] Note that in general, setting up a new modular robot takes around half that time. In this experiment, the most robust solutions required a complete disassembly of an existing configuration before, resulting in a long reconfiguration time.. Both results underline the benefits of multi-objective optimization: Based on the expected disturbances, the required compactness, and the time at hand, an operator can choose the situationally optimal solution for deployment.
The robot could adjust for mobile base positioning errors in all real-world experiments. As expected, the adjustment increases with higher positioning errors. Further, as shown in Figure 5, the absolute errors differ between runs, and for five out of six experiments, the necessary adjustment was below 50% of the maximum adjustment. If a more compact (and thereby less robust) robot along the Pareto-optimal designs needs to be chosen, high positioning errors of the base could likely be overcome by reapproaching the wall and repeating the calibration.
The primary purpose of the conducted experiments was to demonstrate the feasibility of the proposed framework. The gains of the controller were intentionally set to a lower level to prevent excessive force, resulting in reduced precision. This ensures that the robot does not push too far into the wall, even if the vision system fails to accurately detect the marker. Still, as shown in Figure 6, we observed a relative horizontal precision of $1$ cm. In future applications, the system could be enhanced with a more accurate vision system to improve its accuracy. Additionally, implementing a force control mechanism would further robustify the system and enable a more precise and adaptive interaction with the environment.
VI Conclusion
We developed and integrated a holistic framework allowing construction workers to instruct a mobile modular robot arm for building construction. In contrast to existing solutions, the modularity of hardware and software components and multi-objective optimization of robot morphology enables tailoring the proposed approach to specific needs in situ. By integrating BIM information with optimization and control frameworks from the robotics domain, we have created a holistic framework that requires minimal intervention and robotics expertise for deployment. Further, our approach is not limited to geometric data of the used BIM models; future work could incorporate information about material properties to enable automation even for dynamically more challenging tasks, such as impact drilling.
Acknowledgment
The authors gratefully acknowledge financial support by the Horizon 2020 EU Framework Project CONCERT under grant 101016007 and by the Deutsche Forschungsgemeinschaft (German Research Foundation) under grant number AL 1185/31-1. We also thank Matthias Mayer for valuable feedback and discussions and Julius Emig for the help on the experimental setup.